### 17 Mar 2026 v1

# On the norms of the multiplication operators between weighted Bergman spaces

Abstract.
In this paper, we investigate the norms of multiplication operators acting between weighted Bergman spaces. First, we provide a proof for a norm estimate claim previously announced in our recent paper [10]. Second, we establish a sharp norm estimate for certain special multiplication operators between weighted Bergman spaces, a result that appears to be new in the literature.
Key words and phrases:
Multiplication operator, Schwarzian derivative, univalent rational function, norm of Multiplication operator, Koebe function.
2020 Mathematics Subject Classification:
47A30, 47B38
The author was supported by National Natural Science Foundation of China (Grant Nos. 11501157).
1. Introduction and main results
Let $\Delta=\{z:|z|<1\}$ denote the unit disk in the complex plane $\mathbb{C}$. For an open domain $\Omega$ in $\mathbb{C}$, we use $\mathcal{A}(\Omega)$ to denote the class of all analytic functions in $\Omega$. For $1<\gamma<\infty$, we define the weighted Bergman space ${\bf A}_{\gamma}^{2}={\bf A}_{\gamma}^{2}(\Delta)$ as
${\bf A}_{\gamma}^{2}(\Delta)=\{\phi\in\mathcal{A}(\Delta):\|\phi\|_{\gamma}^{2}:=(\gamma+1)\iint_{\Delta}|\phi(z)|^{2}(1-|z|^{2})^{\gamma}\frac{dxdy}{\pi}<\infty\}.$
Let $g\in\mathcal{A}(\Delta)$. The multiplication operator $M_{g}$, induced by $g$, is defined by
$M_{g}(\phi)(z):=g(z)\phi(z),\,\,\phi\in\mathcal{A}(\Delta).$
For $\gamma\in\mathbb{R}$, we define the growth space $\mathcal{B}_{\gamma}=\mathcal{B}_{\gamma}(\Delta)$ as
$\mathcal{B}_{\gamma}(\Delta)=\left\{\phi\in\mathcal{A}(\Delta):\|\phi\|_{\mathcal{B}_{\gamma}}:=\sup_{z\in\Delta}|\phi(z)|(1-|z|^{2})^{\gamma}<\infty\right\},$
which is a Banach space with the norm $\|\cdot\|_{\mathcal{B}_{\gamma}}$ defined above. It is well known that
Theorem 1.1.
Let $\beta>\alpha>-1$. The multiplication operator $M_{g}$ is bounded from ${\bf A}_{\alpha}^{2}$ to ${\bf A}_{\beta}^{2}$ if and only if $g$ belongs to $\mathcal{B}_{\gamma}$ with $\gamma=(\beta-\alpha)/2.$
Remark 1.2.
See [12] for a proof of this result.
For $\beta>\alpha>-1$, when $M_{g}$ is bounded from ${\bf A}_{\alpha}^{2}$ to ${\bf A}_{\beta}^{2}$, we say $g$ is a multiplier from ${\bf A}_{\alpha}^{2}$ to ${\bf A}_{\beta}^{2}$. The class of all multipliers, denoted by $\mathbf{M}_{\alpha,\beta}$, is a Banach space with the following multiplier norm
$\|g\|_{{\bf{M}}_{\alpha,\beta}}:=\sup_{\begin{subarray}{c}\phi\in{\bf A}^{2}_{\alpha}\\ \phi\neq 0\end{subarray}}\frac{\|g\phi\|_{\beta}}{\|\phi\|_{\alpha}}.$
In the paper [11], Shimorin studied a special multiplication operator from ${\bf A}_{\alpha}^{2}$ to ${\bf A}_{\alpha+4}^{2}$, which is induced by the Schwarzian derivative of a univalent function. To recall the results of [11], we first introduce some notations and definitions. We let $\mathcal{S}$ be the class of all univalent functions $f$ in $\Delta$ with $f(0)=f^{\prime}(0)-1=0$.
For a locally univalent function $f$ in an open domain $\Omega$ of $\mathbb{C}$, the Schwarzian derivative $S(f)$ of $f$ which is defined by
$S_{f}(z)=\bigg[\frac{f^{\prime\prime}(z)}{f^{\prime}(z)}\bigg]^{\prime}-\frac{1}{2}\left[\frac{f^{\prime\prime}(z)}{f^{\prime}(z)}\right]^{2},\,z\in\Omega.$
Here, $N(f)(z):=f^{\prime\prime}(z)/f^{\prime}(z)$ is called the Pre-Schwarzian derivative of $f$. The following multiplication operator has been considered by Shimorin in [11].
$M_{S(f)}(\phi)(z):=S(f)(z)\phi(z),\,\phi\in\mathcal{A}(\Delta).$
It is well known that
$|S(f)|(1-|z|^{2})^{2}\leq 6,$
for all $f\in\mathcal{S}$, and hence $S(f)$ belongs to $\mathcal{B}_{2}$, see [3]. Then $S(f)$ is a multiplier from ${\bf A}_{\alpha}^{2}$ to ${\bf A}_{\alpha+4}^{2}$ for any $f\in\mathcal{S}$. Let $t\in\mathbb{R}$. For $f\in\mathcal{S}$, the integral means spectrum $\beta_{f}(t)$ is defined by
(1.1) $\beta_{f}(t):=\limsup_{r\rightarrow 1^{-}}\frac{\log\int_{-\pi}^{\pi}|f^{\prime}(re^{i\theta})|^{t}|d\theta}{|\log(1-r)|}.$
The universal integral means spectrum $B(t)$ is defined by
$B(t)=\sup\limits_{f\in\mathcal{S}}\beta_{f}(t).$
It was proved in [11] that
Theorem 1.3.
Let $\alpha>-1$. Then
$\|S(f)\|_{{\bf{M}}_{\alpha,\alpha+4}}\leq 36\frac{\alpha+3}{\alpha+1},$
holds for all $f\in\mathcal{U}$.
By using this result, Shimorin obtained better bounds for the universal integral means spectrum $B(t)$ for t=1t=-1 and t=2t=-2. B(2)=1B(-2)=1 is known as the famous Brennan conjecture, see [5]. Later, following this line of thought, Hedenmalm and Shimorin established in [7] the current best upper bound estimate about B(2)B(-2) and other universal integral means spectra. For more results about the Brennan conjecture and universal integral means spectrum, see [2], [6].
For simplicity, in the sequel, for a univalent function $f\in\mathcal{S}$, we will write
$\mathbb{M}_{f}(\alpha):=\|S(f)\|_{{\bf M}_{\alpha,\alpha+4}}.$
For a multiplier $g$ from ${\bf A}_{\alpha}^{2}$ to ${\bf A}_{\beta}^{2}$ with $\beta>\alpha>-1$, we will write
$\mathbb{M}_{g}(\alpha,\beta)=\mathbb{M}_{g(z)}(\alpha,\beta):=\|g\|_{{\bf M}_{\alpha,\beta}}.$
There is a close connection between the norm of the multiplication operator ${M}_{S(f)}$ and the Brennan conjecture. It has been shown in [11] that
Proposition 1.4.
Let $\alpha>-1$. We have
$\mathbb{M}_{\kappa}^{2}(\alpha)=\frac{36(\alpha+3)(\alpha+5)}{(\alpha+2)(\alpha+4)},$
Here, $\kappa$ is the well-known Koebe function, which is defined by
$\kappa(z):=\frac{z}{(1-z)^{2}},z\in\Delta.$
If some univalent function $f\in\mathcal{S}$ satisfies that
$\mathbb{M}_{f}^{2}(\alpha)\leq\mathbb{M}_{\kappa}^{2}(\alpha)=\frac{36(\alpha+3)(\alpha+5)}{(\alpha+2)(\alpha+4)},$
for any $\alpha>0$. Then $\beta_{f}(-2)\leq 1$, that is to say, the Brennan conjecture is true for the function $f$.
Let $\mathcal{U}_{R}$ be the class of all univalent rational functions $f$ with $f\in\mathcal{S}$. In the recent paper [10], we have proved that if $R\in\mathcal{U}_{R}$ has at least one critical point on the unit circle $\mathbb{T}$, i.e., there is at least one point $z_{0}$ on $\mathbb{T}$ such that $R^{\prime}(z_{0})=0$, then $\beta_{R}(-2)=1$. On the other hand, by checking carefully the arguments in the proof of Proposition $8$ in [11, Page 1632], we see that, if there is a constant $\varepsilon>0$ such that
$\mathbb{M}_{f}(\alpha)+\mathbb{M}_{f}(\alpha+2)\leq\mathbb{M}_{\kappa}(\alpha)+\mathbb{M}_{\kappa}(\alpha+2),$
for $\alpha\in(0,\varepsilon)$, then $\beta_{f}({-2})\leq 1$. Moreover, for $R\in\mathcal{U}_{R}$ that have at least one critical point on $\mathbb{T}$, if there are two small constants $\epsilon>0,\eta>0$ such that $\mathbb{M}_{R}(\alpha)<\mathbb{M}_{\kappa}(\alpha)-\eta$ for $\alpha\in(-\epsilon,\epsilon)\cup(2-\epsilon,2+\epsilon)$, then we will obtain that $\beta_{R}(-2)<1$, which is a contradiction. This leads us to conjecture the following result. The first aim of this paper is to give a proof for the following theorem.
Theorem 1.5.
For each $\alpha>-1$, we have $\mathbb{M}_{R}(\alpha)\geq\mathbb{M}_{\kappa}(\alpha)$ for all $R\in\mathcal{U}_{R}$ that have at least one critical point on $\mathbb{T}.$
In view of Theorem 1.1, two natural questions are the following
Question 1.6.
For a multiplier $g$ with $\|g\|_{{\mathcal{B}}_{(\beta-\alpha)/2}}=1$, what is its exact multiplier norm?
Question 1.7.
What is the exact value of
$\mathbb{M}(\alpha,\beta):=\sup_{\|g\|_{{\mathcal{B}}_{(\beta-\alpha)/2}}=1}\|g\|_{{\bf M}_{\alpha,\beta}}?$
We will partially answer these two questions. We shall prove that
Theorem 1.8.
Let $\alpha>-1$ and $\beta-\alpha>0$. For $g_{0}(z)=(1-z^{2})^{-\frac{\beta-\alpha}{2}}$, we have
$\mathbb{M}_{g_{0}}^{2}(\alpha,\beta)=\frac{1}{2^{\beta-\alpha}}\frac{\Gamma(\beta+2)}{\Gamma(\alpha+2)}\Big[\frac{\Gamma(1+\alpha/2)}{\Gamma(1+\beta/2)}\Big]^{2}.$
Here $\Gamma$ is the usual Gamma function, see [1].
Remark 1.9.
Note that
$S(\kappa)(z)=-\frac{6}{(1-z^{2})^{2}}.$
Then the first part of Proposition 1.4 follows by taking $\beta=\alpha+4$ in Theorem 1.8. To the best of our knowledge, this result is novel.
From Theorem 1.8, we obtain that
Corollary 1.10.
Let $\beta>\alpha>-1$. Then
$\mathbb{M}^{2}(\alpha,\beta)\geq\frac{1}{2^{\beta-\alpha}}\frac{\Gamma(\beta+1)}{\Gamma(\alpha+1)}\Big[\frac{\Gamma(1+\alpha/2)}{\Gamma(1+\beta/2)}\Big]^{2}.$
The paper is organized as follows. In the next section, we will establish some auxiliary results which will be used in our proof of main results in the paper. We will prove Theorem 1.5 and 1.8 in Section 3 and 4, respectively.
2. Some Auxiliary Results
In this section, we present some auxiliary results which are needed in our proof of main theorems of this paper. First, we need the following integral representation of weighted Bergman spaces. Let $\mathbb{C}_{+}=\{w=u+iv:v>0\}$ be the upper half-plane and $\mathbb{R}_{+}=(0,\infty)$ be the set of positive real numbers. For $\alpha>-1$, we define the weighted Bergman space on the upper half-plane $\bf{A}_{\alpha}^{2}(\mathbb{C}_{+})$ as
${\bf{A}}_{\gamma}^{2}(\mathbb{C}_{+})=\left\{h\in\mathcal{A}(\mathbb{C}_{+}):\|h\|_{\alpha}^{2}=\iint_{\mathbb{C}_{+}}|h(u+iv)|^{2}v^{\alpha}\frac{dudv}{\pi}<\infty\right\}.$
The following lemma will be useful.
Lemma 2.1.
Let $\alpha>-1$. Then $h$ belongs to ${\bf{A}}_{\alpha}^{2}(\mathbb{C}_{+})$ if and only if there is a function $\tilde{h}\in L_{\alpha}^{2}(\mathbb{R}_{+})$ such that
(2.1) $h(w)=\int_{0}^{\infty}\tilde{h}(t)e^{iwt}\,dt,\quad w\in\mathbb{C}_{+},$
and we have the following norm identity
$\|h\|_{\alpha}^{2}=\frac{\Gamma(\alpha+1)}{2^{\alpha}}\int_{0}^{\infty}|\tilde{h}(t)|^{2}\frac{dt}{t^{\alpha+1}}.$
Here $L_{\alpha}^{2}(\mathbb{R}_{+})$ denotes the weighted Lebesgue space consisting of all measurable functions $\tilde{h}$ with norm
$\|\tilde{h}\|_{\alpha}^{2}=\int_{0}^{\infty}|\tilde{h}(t)|^{2}\frac{dt}{t^{\alpha+1}}.$
Remark 2.2.
${\bf{A}}_{\alpha}^{2}(\mathbb{C}_{+})$ is isometrically isomorphic to the space $L_{\alpha}^{2}(\mathbb{R}_{+})$ under the Fourier-Laplace type transform (2.1). See [4] for a proof of this lemma.
We will need the following estimates.
Lemma 2.3.
Let $\alpha>-1$. If $2\lambda>\alpha+2$, then
(2.2) $\|{(1-rz^{2})^{-\lambda}}\|_{\alpha}^{2}=\frac{\Gamma(\alpha+2)\Gamma(2\lambda-\alpha-2)}{2^{\alpha+1}[\Gamma(\lambda)]^{2}}\cdot\frac{1+{o}(1)}{(1-r^{2})^{2\lambda-\alpha-2}},\,{\textup{as}}\,\,r\rightarrow 1^{-}.$
and
(2.3) $\|{(1-rz)^{-\lambda}}\|_{\alpha}^{2}=\frac{\Gamma(\alpha+2)\Gamma(2\lambda-\alpha-2)}{[\Gamma(\lambda)]^{2}}\cdot\frac{1+{o}(1)}{(1-r^{2})^{2\lambda-\alpha-2}},\,{\textup{as}}\,\,r\rightarrow 1^{-}.$
If $0<2\lambda<\alpha+2$, then there is a constant $M>0$ such that for all $r\in(0,1)$,
(2.4) $\|{(1-rz)^{-\lambda}}\|_{\alpha}^{2}\leq M.$
Here and in the sequel, o(1){o}(1) denotes a quantity depending on $r$ such that ${o}(1)\rightarrow 0$ as $r\rightarrow 1^{-}$, and which may vary from line to line.
Proof.
First, for $\gamma>0$, we have
(2.5) $\frac{1}{(1-z)^{\gamma}}=\sum_{n=0}^{\infty}\frac{\Gamma(n+\gamma)}{n!\Gamma(\gamma)}z^{n},z\in\Delta.$
Also, for any $\phi=\sum_{n=0}^{\infty}a_{n}z^{n}\in\mathbf{A}_{\alpha}^{2}(\Delta)$, we have
(2.6) $\|\phi\|_{\alpha}^{2}=\sum_{n=0}^{\infty}\frac{n!\Gamma(\alpha+2)}{\Gamma(n+\alpha+2)}|a_{n}|^{2}.$
Let $r\in(0,1)$ and $2\lambda>\alpha+2$. Then, from (2.5) and (2.6), we obtain that
(2.7) $\|{(1-rz^{2})^{-\lambda}}\|_{\alpha}^{2}=\sum_{n=0}^{\infty}\frac{(2n)!\Gamma(\alpha+2)}{\Gamma(2n+\alpha+2)}\Big|\frac{\Gamma(n+\lambda)}{n!\Gamma(\lambda)}\Big|^{2}r^{2n},$
and
(2.8) $\|{(1-rz)^{-\lambda}}\|_{\alpha}^{2}=\sum_{n=0}^{\infty}\frac{n!\Gamma(\alpha+2)}{\Gamma(n+\alpha+2)}\Big|\frac{\Gamma(n+\lambda)}{n!\Gamma(\lambda)}\Big|^{2}r^{2n}.$
By Stirling’s formula, we have
$\frac{\Gamma(n+\lambda)}{n!}=(n+1)^{\lambda-1}[1+\widetilde{o}(1)],\,\,{\textup{as}}\,\,n\rightarrow\infty.$
Here and later, we use $\widetilde{o}(1)$ to denote a sequence that tends to 0, i.e., $o(1)\rightarrow 0,\,{\textup{as}}\,\,n\rightarrow\infty$, which may be different in different places. Then it follows from (2.7) that
(2.9) $\|{(1-rz^{2})^{-\lambda}}\|_{\alpha}^{2}$ $=$ $\sum_{n=0}^{\infty}\frac{\Gamma(\alpha+2)}{[2(n+1)]^{\alpha+1}}[1+\widetilde{o}(1)]\cdot\frac{(n+1)^{2\lambda-2}}{[\Gamma(\lambda)]^{2}}[1+\widetilde{o}(1)]r^{2n}$
$=$ $\frac{1}{2^{\alpha+1}}\frac{\Gamma(\alpha+2)}{[\Gamma(\lambda)]^{2}}\sum_{n=0}^{\infty}{(n+1)^{2\lambda-\alpha-3}}[1+\widetilde{o}(1)]r^{2n}.$
Similarly, from 2.8, we have
(2.10) $\|{(1-rz)^{-\lambda}}\|_{\alpha}^{2}=\frac{\Gamma(\alpha+2)}{[\Gamma(\lambda)]^{2}}\sum_{n=0}^{\infty}{(n+1)^{2\lambda-\alpha-3}}[1+\widetilde{o}(1)]r^{2n}.$
When a<1a<-1, we have the well-known fact that there is a constant $C>0$ such that
$\sum_{n=1}^{\infty}n^{a}r^{2n}\leq C,$
for all $r\in(0,1)$. Consequently, when $0<2\lambda<\alpha+2$ so that $2\lambda-\alpha-3<-1$, we see that (2.4) holds.
On the other hand, when $2\lambda>\alpha+2$, we have
(2.11) $[{(1-r^{2})^{2\lambda-\alpha-2}}]\sum_{n=0}^{\infty}{(n+1)^{2\lambda-\alpha-3}}[1+\widetilde{o}(1)]r^{2n}$
$=[{(1-r^{2})^{2\lambda-\alpha-2}}]\sum_{n=0}^{\infty}\frac{\Gamma(n+2\lambda-\alpha-2)}{n!\Gamma(2\lambda-\alpha-2)}[1+\widetilde{o}(1)]r^{2n}$
$=1+[{(1-r^{2})^{2\lambda-\alpha-2}}]\sum_{n=0}^{\infty}\frac{\Gamma(n+2\lambda-\alpha-2)}{n!\Gamma(2\lambda-\alpha-2)}\cdot[\widetilde{o}(1)]\cdot r^{2n}$
Note that $\widetilde{o}(1):=b_{n}\rightarrow 0,\,{\textup{as}}\,\,n\rightarrow\infty$ in (2.9). Then, for any small $\varepsilon>0$, there are two constants $n_{0}\in\mathbb{N}$ and $c_{0}>0$ such that $|b_{n}|\leq\varepsilon$ for $n>n_{0}$ and $|b_{n}|\leq c_{0}$ for all $n\in\mathbb{N}\cup\{0\}$. It follows from (2.5) that
(2.12) $[{(1-r^{2})^{2\lambda-\alpha-2}}]\sum_{n=0}^{\infty}\frac{\Gamma(n+2\lambda-\alpha-2)}{n!\Gamma(2\lambda-\alpha-2)}\cdot[\widetilde{o}(1)]\cdot r^{2n}$
$\leq[{(1-r^{2})^{2\lambda-\alpha-2}}]c_{0}\sum_{n=0}^{n_{0}}\frac{\Gamma(n+2\lambda-\alpha-2)}{n!\Gamma(2\lambda-\alpha-2)}r^{2n}$
$\quad\quad\qquad+\varepsilon[{(1-r^{2})^{2\lambda-\alpha-2}}]\sum_{n={{n_{0}}+1}}^{\infty}\frac{\Gamma(n+2\lambda-\alpha-2)}{n!\Gamma(2\lambda-\alpha-2)}r^{2n}$
$\leq c_{0}[{(1-r^{2})^{2\lambda-\alpha-2}}]\sum_{n=0}^{n_{0}}\frac{\Gamma(n+2\lambda-\alpha-2)}{n!\Gamma(2\lambda-\alpha-2)}+\varepsilon$
Combining (2.11) and (2.12), we obtain that
$[{(1-r^{2})^{2\lambda-\alpha-2}}]\sum_{n=0}^{\infty}\frac{\Gamma(n+2\lambda-\alpha-2)}{n!\Gamma(2\lambda-\alpha-2)}[1+\widetilde{o}(1)]r^{2n}\rightarrow 1,{\textup{as}}\,\,r\rightarrow 1^{-}.$
This implies that
$\sum_{n=0}^{\infty}{(n+1)^{2\lambda-\alpha-3}}[1+\widetilde{o}(1)]r^{2n}=\frac{1+{o}(1)}{(1-r^{2})^{2\lambda-\alpha-2}},\,{\textup{as}}\,\,r\rightarrow 1^{-}.$
Then (2.2) and (2.3) follow from (2.9) and (2.10), respectively. The lemma is proved. ∎
The following key lemma will also be needed.
Lemma 2.4.
Let $\alpha>-1,\beta-\alpha>0$. If $g$ is a multiplier from ${\bf A}_{\alpha}^{2}$ to ${\bf A}_{\beta}^{2}$, then for any $a\in\Delta$,
$\mathbb{M}_{g(az)}(\alpha,\beta)\leq\mathbb{M}_{g}(\alpha,\beta).$
Remark 2.5.
We provide a detailed proof of Lemma 2.4, which also clarifies some imprecise statements in [9, page 2310, lines 17-20].
Proof of Lemma 2.4.
Let $\gamma=(\beta-\alpha)/2>0$.We first show that if $g$ is a multiplier from ${\bf A}_{\alpha}^{2}$ to ${\bf A}_{\beta}^{2}$, then the map ${\dagger}:a\mapsto g(az)$ is analytic from the unit disk into $\mathcal{B}_{\gamma}$. Since $g\in\mathcal{B}_{\gamma}$, then for a fixed $a\in\Delta$ and any $z\in\Delta$, we have $1-|z|^{2}\leq 1-|az|^{2}$ so that
$|g(az)|(1-|z|^{2})^{\gamma}=|g(az)|(1-|az|^{2})^{\gamma}\Big[\frac{1-|z|^{2}}{1-|az|^{2}}\Big]^{\gamma}\leq\|g\|_{\mathcal{B}_{\gamma}}.$
This yields that $g(az)\in\mathcal{B}_{\gamma}$ and the mapping $\dagger$} is well-defined.
We will show that $\dagger$} is complex differentiable at an arbitrary point $a_{0}\in\Delta$. Choose $r_{0}>0$ such that $|a_{0}|+r_{0}<1$. Then for any complex $\tau$ with $|\tau|<r_{0}$, we have $a_{0}+\tau\in\Delta$, and for any $z\in\Delta$, $t\in[0,1]$,
$|(a_{0}+t\tau)z|\leq|a_{0}||z|+t|\tau||z|\leq|a_{0}|+|\tau|<|a_{0}|+r_{0}<1.$
For a fixed $z\in\Delta$, when $|\tau|<r_{0}$, we consider the function
$\Psi_{z,\tau}(t)=g((a_{0}+t\tau)z)=g(a_{0}z+t\tau z),\,\,t\in[0,1].$
Since $g$ is analytic, $\psi_{z,h}$ is continuously differentiable on $[0,1]$, and
$\Psi_{z,\tau}^{\prime}(t)=\tau z\,g^{\prime}(a_{0}z+t\tau z).$
Then, by the fundamental theorem of calculus, we have
$\Psi_{z,\tau}(1)-\psi_{z,\tau}(0)=\int_{0}^{1}\Psi_{z,\tau}^{\prime}(t)\,dt.$
That is
$g((a_{0}+\tau)z)-g(a_{0}z)=\tau z\int_{0}^{1}g^{\prime}(a_{0}z+t\tau z)\,dt.$
It follows that for $|\tau|<r_{0}$,
(2.13) $\mathbf{Q}_{\tau}(z):=\frac{g((a_{0}+\tau)z)-g(a_{0}z)}{\tau}=z\int_{0}^{1}g^{\prime}(a_{0}z+t\tau z)\,dt.$
We define $\widetilde{g}(z)=zg^{\prime}(a_{0}z)$. Note that $g^{\prime}$ is continuous on the closed disk $\overline{\Delta}_{r_{1}}$ centered at the origin with radius $r_{1}=\frac{|a_{0}|+r_{0}+1}{2}<1$. Then it is easy to see that $\widetilde{g}\in\mathcal{B}_{\gamma}$.
From (2.13), for $|\tau|<1$, we have
(2.14) $\mathbf{Q}_{\tau}(z)-\widetilde{g}(z)=z\int_{0}^{1}\bigl[g^{\prime}(a_{0}z+t\tau z)-g^{\prime}(a_{0}z)\bigr]\,dt.$
Since the derivative $g^{\prime}$ is uniformly continuous on the closed disk $\overline{\Delta}_{r_{1}}$. For any given $\varepsilon>0$, there exists a $r_{2}>0$ such that for all $w_{1},w_{2}\in\overline{\Delta}_{r_{1}}$ with $|w_{1}-w_{2}|<r_{2}$, we have $|g^{\prime}(w_{1})-g^{\prime}(w_{2})|<\varepsilon$.
Choose $\delta=\min\{r_{0},r_{2}\}$. Note that for any $\tau$ with $|\tau|<\delta$ and for any $z\in\mathbb{D}$ and $t\in[0,1]$, the points $a_{0}z$ and $a_{0}z+t\tau z$ lie in $\overline{\Delta}_{r_{0}}$ because
$|a_{0}z|\leq|a_{0}|<\frac{|a_{0}|+1}{2}<r_{1},$
$|a_{0}z+t\tau z|\leq|a_{0}|+|\tau|\leq|a_{0}|+r_{0}<r_{1}.$
Furthermore,
$|(a_{0}z+t\tau z)-a_{0}z|=|t\tau z|\leq|\tau|<r_{2}.$
Thus the condition for uniform continuity is satisfied and it holds that for $|\tau|<\delta$,
$|g^{\prime}(a_{0}z+t\tau z)-g^{\prime}(a_{0}z)|<\varepsilon,$
for any $t\in[0,1]$, and $z\in\Delta$. It follows from (2.14) that for $|\tau|<\delta$,
$|\mathbf{Q}_{\tau}(z)-\widetilde{g}(z)|\leq|z|\int_{0}^{1}\varepsilon\,dt=\varepsilon|z|.$
Consequently, we obtain that
$\|\mathbf{Q}_{\tau}-\widetilde{g}\|_{\mathcal{B}_{\gamma}}=\sup_{|z|<1}|\mathbf{Q}_{\tau}(z)-\widetilde{g}(z)|(1-|z|^{2})^{\gamma}\leq\varepsilon\sup_{|z|<1}|z|(1-|z|^{2})^{\gamma}\leq\varepsilon.$
Since $\varepsilon$ was arbitrary, we conclude that
$\lim_{\tau\to 0}\|\mathbf{Q}_{\tau}-\widetilde{g}\|_{\mathcal{B}_{\gamma}}=0.$
This means that the mapping ${\dagger}:a\mapsto g(az)$ is analytic from the unit disk to $\mathcal{B}_{\gamma}$.
Now, for a fixed $\phi\in\mathbf{A}_{\alpha}^{2}$, by using the above notations and arguments, we see that $\widetilde{g}\phi$ belongs to $\mathbf{A}_{\beta}^{2}$ and
(2.15) $\|\mathbf{Q}_{\tau}\phi-\widetilde{g}\phi\|_{\beta}\leq\|\mathbf{Q}_{\tau}-\widetilde{g}\|_{\mathcal{B}_{(\beta-\alpha)/2}}\|\phi\|_{\alpha},$
when $\tau$ is close enough to 0. It follows that
$\lim_{\tau\to 0}\|\mathbf{Q}_{\tau}\phi-\widetilde{g}\phi\|_{\beta}=0.$
This proves that
Sub-lemma 2.6.
For a fixed $\phi\in\mathbf{A}_{\alpha}^{2}$, the mapping ${\ddagger}:a\mapsto g(az)\phi(z)$ is analytic from the unit disk to $\mathbf{A}_{\beta}^{2}$.
Next we shall prove that
Sub-lemma 2.7.
For a fixed $\phi\in\mathbf{A}_{\alpha}^{2}$,
$\lim_{\Delta\ni a\to e^{i\theta}}\|g(az)\phi(z)-g(e^{i\theta}z)\phi(z)\|_{\beta}=0,$
for any $\theta\in[0,2\pi)$.
We write $g_{a}(z)=g(az)$. As noted above, we obtain that $\|g_{a}\|_{\mathcal{B}_{\beta-\alpha}}\leq\|g\|_{\mathcal{B}_{\beta-\alpha}}$ so that both $g_{a}\phi$ and $g\phi$ belong to $\mathbf{A}_{\beta}^{2}$. Then from
$\|g(az)\phi(z)-g(e^{i\theta}z)\phi(z)\|_{\beta}^{2}\leq 2\|g_{a}(z)\phi(z)\|_{\beta}^{2}+2\|g(e^{i\theta}z)\phi(z)\|_{\beta}^{2}\leq 4\|g\|_{\mathcal{B}_{(\beta-\alpha)/2}}^{2}\|\phi\|_{\alpha}^{2},$
we see that, for any given $\varepsilon>0$, there is a constant $r_{1}\in(0,1)$ such that
(2.16) ${\bf Int_{1}}:=\frac{\beta+1}{\pi}\int_{|z|>r_{0}}|g(az)\phi(z)-g(e^{i\theta}z)\phi(z)|^{2}(1-|z|^{2})^{\beta}dxdy<\varepsilon.$
We proceed to consider the following integral
(2.17) ${\bf Int_{2}}:=\frac{\beta+1}{\pi}\int_{|z|\leq r_{0}}|g(az)\phi(z)-g(e^{i\theta}z)\phi(z)|^{2}(1-|z|^{2})^{\beta}dxdy.$
Note that, on the closed disk $\{|z|\leq r_{0}\}$, $g(az)$ converges uniformly to $g(e^{i\theta}z)$ as $\Delta\ni a\to e^{i\theta}$, i.e.,
$\lim_{\Delta\ni a\to e^{i\theta}}\sup_{|z|\leq r_{0}}|g(az)-g(e^{i\theta}z)|=0.$
Hence for given $\varepsilon>0$, there is a constant $\delta>0$ such that
$\sup_{|z|\leq r_{0}}|g(az)-g(e^{i\theta}z)|(1-|z|^{2})^{(\beta-\alpha)/2}<\sqrt{\varepsilon},$
for all $a\in\Delta$ with $|a-e^{i\theta}|<\delta$. Consequently, when $a\in\Delta$ with $|a-e^{i\theta}|<\delta$, we have
(2.18) ${\bf Int_{2}}\leq{\varepsilon}\frac{\beta+1}{\pi}\int_{|z|\leq r_{0}}|\phi(z)|^{2}(1-|z|^{2})^{\alpha}dxdy\leq{\varepsilon}\frac{\beta+1}{(\alpha+1)\pi}\|\phi\|_{\alpha}^{2}.$
From (2.16) and (2.18), we conclude that
$\lim_{\Delta\ni a\to e^{i\theta}}\|g(az)\phi(z)-g(e^{i\theta}z)\phi(z)\|_{\beta}^{2}=0,$
for any $\theta\in[0,2\pi)$. This proves Sub-Lemma 2.7.
Finally, for any $\theta\in\mathbb{R}$, it is easy to check that $\|\phi(e^{i\theta}z)\|_{\alpha}=\|\phi(z)\|_{\alpha}$ for any $\phi\in\mathbf{A}_{\alpha}^{2}$. Then we can obtain by a change of variables that
$\mathbb{M}_{g(e^{i\theta}z)}(\alpha,\beta)=\mathbb{M}_{g}(\alpha,\beta),$
for any $\theta\in[0,2\pi)$. On the other hand, by means of Sub-Lemma 2.6 and 2.7, we know that, for a fixed $\phi\in\mathbf{A}_{\alpha}^{2}$, the mapping
${\ddagger}:a\mapsto g(az)\phi(z),$
is analytic from the unit disk $\Delta$ to $\mathbf{A}_{\beta}^{2}$ and continuous on the unit circle. Then $\|g(az)\phi(z)\|_{\beta}$ is subharmonic in $\Delta$, see [8]. It follows from the maximum modulus principle that
$\mathbb{M}_{g_{a}}(\alpha,\beta)\leq\mathbb{M}_{g}(\alpha,\beta),$
for any $a\in\Delta$. This completes the proof of Lemma 2.4. ∎
We also need the following weighted Hardy inequality. For completeness, we include a proof here for this result.
Lemma 2.8.
Let $p>1,r>0$, a<p1a<p-1. Let $f$ be a measurable function on $\mathbb{R}_{+}$ such that $\int_{0}^{\infty}x^{a}|f(x)|^{p}\,dx<\infty$. Then we have
(2.19) $\int_{0}^{\infty}x^{a}\Bigl|\frac{f_{r}(x)}{x^{r}}\Bigr|^{p}dx\leq\Bigl[\frac{\Gamma\bigl(1-\frac{a+1}{p}\bigr)}{\Gamma\bigl(r+1-\frac{a+1}{p}\bigr)}\Bigr]^{p}\int_{0}^{\infty}x^{a}|f(x)|^{p}\,dx,$
where
$f_{r}(x)=\frac{1}{\Gamma(r)}\int_{0}^{x}(x-t)^{r-1}f(t)\,dt.$
Proof.
For fixed $x>0$, let $t=xu$ with $u\in(0,1)$. Then
$\frac{f_{r}(x)}{x^{r}}=\frac{1}{\Gamma(r)}\int_{0}^{1}(1-u)^{r-1}f(xu)\,du.$
Let $K(u)=\frac{(1-u)^{r-1}}{\Gamma(r)}$. By the Minkowski integral inequality,
$\left(\int_{0}^{\infty}x^{a}\left|\frac{f_{r}(x)}{x^{r}}\right|^{p}dx\right)^{1/p}\leq\int_{0}^{1}K(u)\left(\int_{0}^{\infty}x^{a}|f(xu)|^{p}dx\right)^{1/p}du.$
For fixed $u>0$, let $y=xu$ so that
$\int_{0}^{\infty}x^{a}|f(xu)|^{p}dx=u^{-a-1}\int_{0}^{\infty}y^{a}|f(y)|^{p}dy.$
Consequently,
(2.20) $\left(\int_{0}^{\infty}x^{a}\left|\frac{f_{r}(x)}{x^{r}}\right|^{p}dx\right)^{1/p}\leq\left(\int_{0}^{\infty}y^{a}|f(y)|^{p}dy\right)^{1/p}\int_{0}^{1}K(u)u^{-(a+1)/p}du.$
On the other hand, by the definition of Gamma function, we have
$\int_{0}^{1}K(u)u^{-(a+1)/p}du$ $=$ $\frac{1}{\Gamma(r)}\int_{0}^{1}(1-u)^{r-1}u^{-(a+1)/p}du$
$=$ $\frac{\Gamma\left(1-\frac{a+1}{p}\right)}{\Gamma\left(r+1-\frac{a+1}{p}\right)}.$
Then the inequality (2.19) follows from (2.20) and (2). The lemma is proved. ∎
3. Proof of Theorem 1.5
To prove Theorem 1.5, we first recall some notions and results from [10]. For a number $n\in\mathbb{N}_{0}=\mathbb{N}\cup\{0\}$, we will use the notation $P_{n}(z)$ to denote a polynomial of degree at most $n$. We denote by $\mathcal{R}$ the class of all rational functions $R$ in $\Delta$ with $R(0)=R^{\prime}(0)-1=0$ and which have no poles in $\Delta$. For a subset $\Omega\subset\mathbb{C}$, we say that $R\in\mathcal{R}$ has no zeros, or one zero in $\Omega$, if the equation $R(z)=0$ has no solutions, or exactly one solution in $\Omega$, respectively. We see that a rational function $R\in\mathcal{R}$ can be uniquely represented as
$R(z)=\frac{P_{n}(z)}{P_{m}(z)}=\frac{z+a_{2}z^{2}+\cdots+a_{n}z^{n}}{1+b_{1}z+\cdots+b_{m}z^{m}}.$
Here, $P_{n}$ and $P_{m}$ have no common factors except 1 and $P_{m}(z)$ has no zeros in $\Delta$.
When $R(z)$ belongs to $\mathcal{U}_{R}$, we know from the Koebe distortion theorem that
$|R(z)|\leq\frac{|z|}{(1-|z|)^{2}}.$
In view of this fact, we shall consider the following three three subclasses of $\mathcal{R}$.
Definition 3.1.
$\bullet$} Class $\mathcal{L}_{I}$}. The class $\mathcal{L}_{I}$} consists of all functions $R\in\mathcal{R}$ that can be written as
(3.1) $R(z)=\frac{P_{n}(z)}{P_{m}(z)}.$
Here, $P_{n}$ and $P_{m}$ have no common factors except $1$, and $P_{m}(z)$ has no zeros in $\overline{\Delta}=\Delta\cup\mathbb{T}$. It is easy to see that $\mathcal{U}_{P}\subset\mathcal{L}_{I}.$
$\bullet$} Class $\mathcal{L}_{II}$}. The class $\mathcal{L}_{II}$} consists of all functions $R\in\mathcal{R}$ that can be written as
(3.2) $R(z)=\frac{P_{n}(z)}{\prod\limits_{j=1}^{l}(z-e^{i\theta_{j}})P_{m}(z)}.$
Here, $P_{m}$ has no zeros in $\overline{\Delta}$. $l\in\mathbb{N},\theta_{j}\in(-\pi,\pi]$, $\theta_{j_{1}}\neq\theta_{j_{2}}$ for $j_{1}\neq j_{2}$, the numerator and denominator of $R$ have no common factors except $1$.
$\bullet$} Class $\mathcal{L}_{III}$}. The class $\mathcal{L}_{III}$} consists of all functions $R\in\mathcal{R}$ that can be written as
(3.3) $R(z)=\frac{P_{n}(z)}{\prod\limits_{j=1}^{l}(z-e^{i\theta_{j}})^{2}\prod\limits_{k=1}^{t}(z-e^{i\widetilde{\theta}_{k}})P_{m}(z)}.$
Here, $P_{m}$ has no zeros in $\overline{\Delta}$. $l\in\mathbb{N},\theta_{j}\in(-\pi,\pi]$, $\theta_{j_{1}}\neq\theta_{j_{2}}$ for $j_{1}\neq j_{2}$, $t\in\mathbb{N}\cup\{0\}$, $\widetilde{\theta}_{k}\in(-\pi,\pi]$, $\widetilde{\theta}_{k_{1}}\neq\widetilde{\theta}_{k_{2}}$ for $k_{1}\neq k_{2}$. Moreover, ${\theta}_{j}\neq\widetilde{\theta}_{k}$ for all $j,k$. The numerator and denominator of $R$ have no common factors except $1$. In particular, the case $t=0$ means that the denominator of $R$ is $\prod\limits_{j=1}^{l}(z-e^{i\theta_{j}})^{2}P_{m}(z).$
Remark 3.2.
We note that $\mathcal{U}_{R}\subset\mathcal{L}_{I}\cup{\mathcal{L}_{II}}\cup{\mathcal{L}_{III}}$.
On the other hand, when $R$ belongs to $\mathcal{U}_{R}$, we know that
$|R^{\prime}(z)|\geq\frac{1-|z|}{(1+|z|)^{3}}.$
Hence, if $R\in\mathcal{U}_{R}$ has at least one critical point on $\mathbb{T}$, then $R^{\prime}(z)$ can be written as the following unique form.
$R^{\prime}(z)=\frac{[(z-z_{1})(z-z_{2})\cdots(z-z_{s})]P_{\bf n}(z)}{P_{\bf m}(z)}.$
Here, $s\in\mathbb{N}$, and $z_{j},j=1,2,\cdots,s$ are all distinct critical points of $R$ on $\mathbb{T}$. The numerator and denominator of $R^{\prime}$ have no common factors except $1$. $P_{\bf n}(z)$ has no zeros on $\mathbb{T}$, and $P_{\bf m}(z)$ has no zeros in $\Delta$.
Now, we introduce another subclass of $\mathcal{R}$ as follows.
Definition 3.3.
For a function $R\in\mathcal{R}$, we say it belongs to the class $\mathcal{R}_{O}$, if $R^{\prime}$ can be written uniquely in the following form:
(3.4) $R^{\prime}(z)=\frac{[(z-z_{1})(z-z_{2})\cdots(z-z_{s})]P_{\bf n}(z)}{P_{\bf m}(z)}:=\frac{\Pi(z)P_{\bf n}(z)}{P_{\bf m}(z)}.$
Here $s\in\mathbb{N}\cup\{0\}$. When $s\geq 1$, $z_{j},j=1,2,\cdots,s$ are all distinct critical points on $\mathbb{T}$ of $R$. The numerator and denominator of $R^{\prime}$ have no common factors except $1$. $P_{\bf n}(z)$ has no zeros on $\mathbb{T}$, and $P_{\bf m}(z)$ has no zeros in $\Delta$. The case $s=0$ means that $\Pi(z)\equiv 1$ and $R$ has no critical points on $\mathbb{T}$. Note that we allow here $P_{\bf n}(z)$ to have zeros in the unit disk.
Remark 3.4.
It is easy to see that $\mathcal{U}_{R}$ is contained in $\mathcal{R}_{O}\cap({\mathcal{L}_{I}}\cup{\mathcal{L}_{II}}\cup{\mathcal{L}_{III}}).$
For $R\in\mathcal{R}_{O}$, we let, as in (3.4),
$R^{\prime}(z)=\frac{[(z-z_{1})(z-z_{2})\cdots(z-z_{s})]P_{\bf n}(z)}{P_{\bf m}(z)}.$
Furthermore, (a) If $R$ belongs to the class $\mathcal{R}_{O}\cap\mathcal{L}_{I}$, we let, as in (3.1),
$R(z)=\frac{P_{n}(z)}{P_{m}(z)}.$
Then
(3.5) $R^{\prime}(z)=\frac{P^{\prime}_{n}(z)P_{m}(z)-P_{n}(z)P^{\prime}_{m}(z)}{[P_{m}(z)]^{2}}=\frac{[(z-z_{1})(z-z_{2})\cdots(z-z_{s})]P_{\bf n}(z)}{P_{\bf m}(z)}$
Note that the numerator and denominator of the last rational function in (3.5) have no common factors except $1$ so that $P_{\bf m}(z)$ has no zeros in $\overline{\Delta}.$
(b) If $R$ belongs to the class $\mathcal{R}_{O}\cap\mathcal{L}_{II}$, we let, as in (3.2),
$R(z)=\frac{P_{n}(z)}{\prod\limits_{j=1}^{l}(z-e^{i\theta_{j}})P_{m}(z)}:=\frac{1}{\Pi_{1}}\frac{P_{n}(z)}{P_{m}(z)}.$
Then
$R^{\prime}(z)$ $=$ $\frac{\Pi_{1}[P^{\prime}_{n}(z)P_{m}(z)-P_{n}(z)P^{\prime}_{m}(z)]-\Pi_{1}^{\prime}P_{n}(z)P_{m}(z)}{\Pi_{1}^{2}[P_{m}(z)]^{2}}$
$=$ $\frac{[(z-z_{1})(z-z_{2})\cdots(z-z_{s})]P_{\bf n}(z)}{\Pi_{1}^{2}P_{b}(z)}.$
Here, the numerator and denominator of the last rational function in (3) have no common factors except $1$ so that $P_{\bf m}(z)=\Pi_{1}^{2}P_{b}(z)=\Pi_{1}^{2}P_{{\bf m}-2l}(z)$, and $P_{b}(z)=P_{{\bf m}-2l}(z)$ has no zeros in $\overline{\Delta}$.
(c) If $R$ belongs to the class $\mathcal{R}_{O}\cap\mathcal{L}_{III}$, we let, as in (3.3),
$R(z)=\frac{P_{n}(z)}{\prod\limits_{j=1}^{l}(z-e^{i\theta_{j}})^{2}\prod\limits_{k=1}^{t}(z-e^{i\widetilde{\theta}_{k}})P_{m}(z)}:=\frac{1}{\Pi_{2}}\frac{P_{n}(z)}{P_{m}(z)}.$
Then
$R^{\prime}(z)$ $=$ $\frac{\Pi_{2}[P^{\prime}_{n}(z)P_{m}(z)-P_{n}(z)P^{\prime}_{m}(z)]-\Pi_{2}^{\prime}P_{n}(z)P_{m}(z)}{\Pi_{2}^{2}[P_{m}(z)]^{2}}$
$=$ $\frac{[(z-z_{1})(z-z_{2})\cdots(z-z_{s})]P_{\bf n}(z)}{\prod\limits_{j=1}^{l}(z-e^{i\theta_{j}})^{3}\prod\limits_{k=1}^{t}(z-e^{i\widetilde{\theta}_{k}})^{2}P_{b}(z)}.$
Here, the numerator and denominator of the last rational function in (3) have no common factors except $1$ so that
$P_{\bf m}(z)=\prod\limits_{j=1}^{l}(z-e^{i\theta_{j}})^{3}\prod\limits_{k=1}^{t}(z-e^{i\widetilde{\theta}_{k}})^{2}P_{b}(z)=\prod\limits_{j=1}^{l}(z-e^{i\theta_{j}})^{3}\prod\limits_{k=1}^{t}(z-e^{i\widetilde{\theta}_{k}})^{2}P_{{\bf m}-3l-2t}(z),$
and $P_{b}(z)=P_{{\bf m}-3l-2t}(z)$ has no zeros in $\overline{\Delta}$.
Therefore, based on the above arguments, we next will consider the following subclasses $\mathcal{Q}_{I}$, $\mathcal{Q}_{II}$, $\mathcal{Q}_{III}$ of the class of all locally univalent functions in $\Delta$ for the purpose of the paper.
Definition 3.5.
Let $R$ be a locally univalent function in $\Delta$.
$\bullet$ We say $R$ belongs to the class $\mathcal{Q}_{I}$, if $R^{\prime}$ can be written in the following form
(3.8) $R^{\prime}(z)=\frac{[(z-z_{1})(z-z_{2})\cdots(z-z_{s})]P_{\bf n}(z)}{P_{\bf m}(z)}:=\frac{\Pi(z)P_{\bf n}(z)}{P_{\bf m}(z)}.$
Here $s\in\mathbb{N}\cup\{0\}$. When $s\geq 1$, $z_{j},j=1,2,\cdots,s$ are all distinct critical points on $\mathbb{T}$ of $R$. The numerator and denominator of $R^{\prime}$ have no common factors except $1$. $P_{\bf n}(z)$ and $P_{\bf m}(z)$ have no zeros in $\overline{\Delta}$. The case $s=0$ means that $\Pi(z)\equiv 1$ and $R$ has no critical points on $\mathbb{T}$.
$\bullet$ We say $R$ belongs to the class $\mathcal{Q}_{II}$, if $R^{\prime}$ can be written in the following form
(3.9) $R^{\prime}(z)=\frac{[(z-z_{1})(z-z_{2})\cdots(z-z_{s})]P_{\bf n}(z)}{\prod\limits_{j=1}^{l}(z-e^{i\theta_{j}})^{2}P_{{\bf m}}(z)}.$
Here $s\in\mathbb{N}\cup\{0\},l\in\mathbb{N}$. When $s\geq 1$, $z_{j},j=1,2,\cdots,s$ are all distinct critical points on $\mathbb{T}$ of $f$. The numerator and denominator of $R^{\prime}$ in (3.9) have no common factors except $1$. $P_{\bf n}(z)$ and $P_{\bf m}(z)$ have no zeros in $\overline{\Delta}$. In particular, $R^{\prime}$ has no zeros on $\mathbb{T}$ when $s=0$.
$\bullet$ We say $R$ belongs to the class $\mathcal{Q}_{III}$, if $R^{\prime}$ can be written in the following form
(3.10) $R^{\prime}(z)$ $=$ $\frac{[(z-z_{1})(z-z_{2})\cdots(z-z_{s})]P_{\bf n}(z)}{\prod\limits_{j=1}^{l}(z-e^{i\theta_{j}})^{3}\prod\limits_{k=1}^{t}(z-e^{i\widetilde{\theta}_{k}})^{2}P_{\bf m}(z)}.$
Here $s,t\in\mathbb{N}\cup\{0\},l\in\mathbb{N}$. When $s\geq 1$, $z_{j},j=1,2,\cdots,s$ are all distinct critical points on $\mathbb{T}$ of $f$. The numerator and denominator of $R^{\prime}$ in (3.10) have no common factors except $1$. $P_{\bf n}(z)$ and $P_{\bf m}(z)$ have no zeros in $\overline{\Delta}$. In particular, $R^{\prime}$ has no zeros on $\mathbb{T}$ when $s=0$ and $R^{\prime}$ has no poles of order $2$ on $\mathbb{T}$ when $t=0$.
Remark 3.6.
It is easy to see that $\mathcal{U}_{R}$ is contained in ${\mathcal{Q}_{I}}\cup{\mathcal{Q}_{II}}\cup{\mathcal{Q}_{III}}.$
Now, to prove Theorem 1.5, it is enough to show that following result holds.
Proposition 3.7.
Let $\alpha>-1$. Let $R$ be a locally univalent function in $\Delta$ and belong to $\mathcal{Q}_{I}\cup\mathcal{Q}_{II}\cup\mathcal{Q}_{III}$. If $R$ has at least one critical point on $\mathbb{T}$, then $S(R)$ is a multiplier from ${\bf A}_{\alpha}^{2}$ to ${\bf A}_{\alpha+4}^{2}$ and
$\|S(R)\|_{{\bf M}_{\alpha,\alpha+4}}\geq\mathbb{M}_{\kappa}(\alpha).$
Proof of Proposition 3.7.
Assume that $R$ is a locally univalent function in $\Delta$ with having at least one critical point on $\mathbb{T}$. We first show that $S(R)$ is a multiplier from ${\bf A}_{\alpha}^{2}$ to ${\bf A}_{\alpha+4}^{2}$.
When $R$ belongs to $\mathcal{Q}_{I}$, as in 3.8, we have
$R^{\prime}(z)=\frac{[(z-z_{1})(z-z_{2})\cdots(z-z_{s})]P_{\bf n}(z)}{P_{\bf m}(z)},s\geq 1.$
A direct computation gives that
$N(R)(z)=\sum_{r=1}^{s}\frac{1}{z-z_{r}}+\frac{P^{\prime}_{\bf n}(z)}{P_{\bf n}(z)}-\frac{P^{\prime}_{\bf m}(z)}{P_{\bf m}(z)}.$
It follows that
$N(R)(z)]^{\prime}=-\sum_{r=1}^{s}\frac{1}{(z-z_{r})^{2}}+\Big[\frac{P^{\prime}_{\bf n}(z)}{P_{\bf n}(z)}\Big]^{\prime}-\Big[\frac{P^{\prime}_{\bf m}(z)}{P_{\bf m}(z)}\Big]^{\prime}.$
Then, from the definition of Schwarzian derivative
$S(R)(z)=[N(R)(z)]^{\prime}-\frac{1}{2}[N(R)(z)]^{2},$
and by using the element identity
$(\sum_{m=1}^{n}a_{m})^{2}=\sum_{m=1}^{n}{a_{m}^{2}}+2\sum_{i=1}^{n-1}\sum_{j=i+1}^{n}a_{i}a_{j},\,\,n\geq 2.$
we conclude that $S(R)$ has the following form
(3.11) $S(R)(z)=-\frac{3}{2}\sum_{r=1}^{s}\frac{1}{(z-z_{r})^{2}}+\sum_{i=1}^{s}\frac{A_{r}(z)}{z-z_{s}}+{\bf Anal}(z).$
Here,
(3.12) ${\bf Anal}(z)=\Big[\frac{P^{\prime}_{\bf n}(z)}{P_{\bf n}(z)}\Big]^{\prime}-\Big[\frac{P^{\prime}_{\bf m}(z)}{P_{\bf m}(z)}\Big]^{\prime}-\frac{1}{2}\Big[\frac{P^{\prime}_{\bf n}(z)}{P_{\bf n}(z)}\Big]^{2}-\frac{1}{2}\Big[\frac{P^{\prime}_{\bf m}(z)}{P_{\bf m}(z)}\Big]^{2},$
For each $r\in[1,s]$, $A_{r}(z)$ is analytic in $\overline{\Delta}$.
When $R$ belongs to $\mathcal{Q}_{II}$, as in 3.9, we have
$R^{\prime}(z)=\frac{[(z-z_{1})(z-z_{2})\cdots(z-z_{s})]P_{\bf n}(z)}{\prod\limits_{j=1}^{l}(z-e^{i\theta_{j}})^{2}P_{{\bf m}}(z)},l\geq 1,s\geq 1.$
Then we have
$N(R)(z)=\sum_{r=1}^{s}\frac{1}{z-z_{r}}-2\sum_{j=1}^{l}\frac{1}{z-e^{i\theta_{j}}}+\frac{P^{\prime}_{\bf n}(z)}{P_{\bf n}(z)}-\frac{P^{\prime}_{\bf m}(z)}{P_{\bf m}(z)},$
so that
$N(R)(z)]^{\prime}=-\sum_{r=1}^{s}\frac{1}{(z-z_{r})^{2}}+2\sum_{j=1}^{l}\frac{1}{(z-e^{i\theta_{j}})^{2}}+\Big[\frac{P^{\prime}_{\bf n}(z)}{P_{\bf n}(z)}\Big]^{\prime}-\Big[\frac{P^{\prime}_{\bf m}(z)}{P_{\bf m}(z)}\Big]^{\prime}.$
Then, similarly as above, we have
$S(R)(z)=-\frac{3}{2}\sum_{r=1}^{s}\frac{1}{(z-z_{r})^{2}}+\sum_{r=1}^{s}\frac{B_{r}(z)}{z-z_{r}}$
$\quad\quad\,\,\,\,-2\sum_{j=1}^{l}\frac{1}{(z-e^{i\theta_{j}})^{2}}+\sum_{j=1}^{l}\frac{C_{j}(z)}{z-e^{i\theta_{j}}}+{\bf Anal}(z).$
Here, ${\bf Anal}(z)$ is defined as in (3.12). for each $r\in[1,s]$, $j\in[1,l]$, $B_{i}(z)$ and $C_{j}(z)$ are analytic in $\overline{\Delta}$.
When $R$ belongs to $\mathcal{Q}_{III}$, as in 3.10, we have
$R^{\prime}(z)=\frac{[(z-z_{1})(z-z_{2})\cdots(z-z_{s})]P_{\bf n}(z)}{\prod\limits_{j=1}^{l}(z-e^{i\theta_{j}})^{3}\prod\limits_{k=1}^{t}(z-e^{i\widetilde{\theta}_{k}})^{2}P_{\bf m}(z)},\,s\geq 1,l\geq 1,t\geq 0.$
Then we have
$N(R)(z)=\sum_{r=1}^{s}\frac{1}{z-z_{r}}-3\sum_{j=1}^{l}\frac{1}{z-e^{i\theta_{j}}}-2\sum_{k=1}^{t}\frac{1}{z-e^{i\widetilde{\theta}_{k}}}+\frac{P^{\prime}_{\bf n}(z)}{P_{\bf n}(z)}-\frac{P^{\prime}_{\bf m}(z)}{P_{\bf m}(z)},$
and
$[N(R)(z)]^{\prime}=-\sum_{r=1}^{s}\frac{1}{(z-z_{r})^{2}}+3\sum_{j=1}^{l}\frac{1}{(z-e^{i\theta_{j}})^{2}}$
$\quad\quad\quad\quad+2\sum_{k=1}^{t}\frac{1}{(z-e^{i\widetilde{\theta}_{k}})^{2}}+\Big[\frac{P^{\prime}_{\bf n}(z)}{P_{\bf n}(z)}\Big]^{\prime}-\Big[\frac{P^{\prime}_{\bf m}(z)}{P_{\bf m}(z)}\Big]^{\prime}.$
Consequently, similarly again as above, we have, for $t\geq 1$,
$S(R)(z)=-\frac{3}{2}\sum_{r=1}^{s}\frac{1}{(z-z_{r})^{2}}+\sum_{r=1}^{s}\frac{D_{r}(z)}{z-z_{r}}$
$\quad\quad\,\,\,\,-\frac{3}{2}\sum_{j=1}^{l}\frac{1}{(z-e^{i\theta_{j}})^{2}}+\sum_{j=1}^{l}\frac{E_{j}(z)}{z-e^{i\theta_{j}}}$
$\quad\quad\,\,\,\,-2\sum_{k=1}^{t}\frac{1}{(z-e^{i\widetilde{\theta}_{k}})^{2}}+\sum_{k=1}^{t}\frac{F_{k}(z)}{z-e^{i\widetilde{\theta}_{k}}}+{\bf Anal}(z).$
Here, ${\bf Anal}(z)$ is the same as in (3.12) and $D_{r},E_{j},F_{k}$ are all analytic in $\Delta$ for each $r\in[1,s],j\in[1,l],k\in[1,t]$. Note that when $t=0$, there is no the following terms
$2\sum_{k=1}^{t}\frac{1}{(z-e^{i\widetilde{\theta}_{k}})^{2}}+\sum_{k=1}^{t}\frac{F_{k}(z)}{z-e^{i\theta_{j}}},$
in the right side of 3. Combining (3.11), (3) and (3.11), we can see that $S(R)$ is a multiplier from ${\bf A}_{\alpha}^{2}$ to ${\bf A}_{\alpha+4}^{2}$.
To proceed the proof, we need the following claim.
Claim 3.8.
Let $\alpha>-1$ and let $\alpha+2<2\lambda<\alpha+3$. Let $A(z)$ be analytic in $\overline{\Delta}$.
(1) We define
$\mathbf{T}_{1}:=\frac{\alpha+5}{\pi}\iint_{\Delta}|A(z)||1-rz|^{-2\lambda-3}(1-|z|^{2})^{\alpha+4}dxdy,$
and
$\mathbf{T}_{2}:=\frac{\alpha+5}{\pi}\iint_{\Delta}|A(z)||1-rz|^{-2\lambda-2}(1-|z|^{2})^{\alpha+4}dxdy,$
Then there is a constant $M>0$ such that ${\bf T}_{r}\leq M$ for all $r=1,2$ and all $r\in(0,1)$.
(2) Let $\Theta\in(-2\pi,2\pi)$ be such that $e^{-i\Theta}\neq 1$. We define
$\mathbf{T}_{3}:=\frac{\alpha+5}{\pi}\iint_{\Delta}|A(z)||1-re^{-i\Theta}z|^{-1}|1-rz|^{-2\lambda-2}(1-|z|^{2})^{\alpha+4}dxdy,$
and
$\mathbf{T}_{4}:=\frac{\alpha+5}{\pi}\iint_{\Delta}|A(z)||1-re^{-i\Theta}z|^{-2}|1-rz|^{-2\lambda-2}(1-|z|^{2})^{\alpha+4}dxdy,$
Then there is a constant $M>0$ such that ${\bf T}_{r}\leq M$ for all $i=3,4$ and all $r\in(0,1)$.
Proof.
First, note that there is a constant $M_{1}>0$ such that $|A(z)|\leq M_{1}$ for all $z\in\Delta$ since $A(z)$ be analytic in $\overline{\Delta}$. Then we have
${\bf T}_{1}\leq M_{1}\|(1-rz)^{-\lambda-\frac{3}{2}}\|_{\alpha+4}^{2},\,\,{\bf T}_{2}\leq M_{1}\|(1-rz)^{-\lambda-1}\|_{\alpha+4}^{2}.$
Consequently, by using the third case of Lemma 2.3 and its proof, since $2\lambda+2<2\lambda+3<(\alpha+4)+2$, we see that there is a constant $M>0$ such that ${\bf T}_{r}\leq M$ for all $r=1,2$ and all $r\in(0,1)$. This proves the case (1).
We next only prove the case for ${\bf T}_{4}$, The case for ${\bf T}_{3}$ can be similarly proved. In view of the conditions of the case (2), for some fixed $r_{0}\in(0,1)$, we can take two constants $\eta_{1}>0,\eta_{2}>0$ such that $\overline{\Omega}_{1}\cap\overline{\Omega}_{2}=\emptyset$. Here
$\Omega_{1}=\{z=re^{i\theta}:r_{0}<r<1,\Theta-\eta_{1}<\theta<\Theta+\eta_{1}\},$
$\Omega_{=}\{z=re^{i\theta}:r_{0}<r<1,-\eta_{2}<\theta<\eta_{2}\}.$
We set
$\phi_{r}(z)=\frac{\alpha+5}{\pi}|1-re^{-i\Theta}z|^{-2}|1-rz|^{-2\lambda-2}(1-|z|^{2})^{\alpha+4}.$
Since $|A(z)|$ is bounded in $\Delta$, it is enough to prove that $\iint_{\Delta}\phi_{r}(z)dxdy<M$ for some $M>0$. Note that
$\sup_{r\in(0,1),z\in\Omega_{1}}|1-rz|^{-2\lambda-2}=\max_{z\in{\pi_{1}}}\{|1-z|^{-2\lambda-2}\},$
here, ${\pi}_{1}$ is the sector $\{z=re^{i\theta}:r\in[0,1],\theta\in[-\eta_{1}-\Theta,\eta_{2}+\Theta]$. Since 1z1-z is continuous in the compact set ${\bf\pi}_{1}$ and the point $1$ does not belongs to ${\bf\pi}_{1}$, then we see that there is a constant $C_{1}>0$ such that
$|1-rz|^{-2\lambda-2}\leq C_{1},$
for all $z\in\Omega_{1}$ and $r\in(0,1)$. Similarly, since
$\sup_{r\in(0,1),z\in\Omega_{2}}|1-re^{i\Theta}z|^{-2}=\sup_{r\in(0,1),z\in\Omega_{2}}|e^{-i\Theta}-rz|^{-2}=\max_{z\in{\pi_{2}}}\{|e^{-i\Theta}-z|^{-2}\},$
here, ${\bf\pi}_{2}$ is the sector $\{z=re^{i\theta}:r\in[0,1],\theta\in[-\eta_{2},\eta_{2}]$. As $e^{-i\Theta}-z$ is continuous in the compact set $\pi_{2}$ and the point $e^{-i\Theta}$ is not contained in $\pi_{2}$, hence we know that there is a constant $C_{2}>0$ such that
$|1-re^{i\Theta}z|^{-2}\leq C_{2},$
for all $z\in\Omega_{2}$ and $r\in(0,1)$. Meanwhile, we have
[…]
here, $\pi_{3}$ is the following compact set $\pi_{3}=\overline{\Delta}-s_{1}-s_{2}$, where
$s_{1}:=\{z=re^{i\theta}:0<r\leq 1,-\eta_{1}+\Theta<\theta<\eta_{1}+\Theta\},$
and
$s_{2}:=\{z=re^{i\theta}:0<r\leq 1,-\eta_{2}<\theta<\eta_{2}\}.$
Therefore, from the facts that both the points $1$ and $e^{-i\Theta}$ are not contained in $\pi_{3}$, and $1-z,e^{-i\Theta}-z$ are continuous in $\pi_{3}$, we see that there is a constant $C_{3}>0$ such that
$|1-re^{i\Theta}z|^{-2}|1-rz|^{-2\lambda-2}\leq C_{3},$
for all $z\in\Delta-\Omega_{1}-\Omega_{2}$ and $r\in(0,1)$. Consequently, we obtain that
$\iint_{\Omega_{1}}\phi_{r}(z)dxdy$ $\leq$ $C_{1}\frac{\alpha+5}{\pi}\iint_{\Omega_{1}}|1-re^{-i\Theta}z|^{-2}(1-|z|^{2})^{\alpha+4}dxdy$
$\leq$ $C_{1}\|(1-rz)^{-1}\|_{\alpha+4}^{2},$
and
$\iint_{\Omega_{2}}\phi_{r}(z)dxdy$ $\leq$ $C_{2}\frac{\alpha+5}{\pi}\iint_{\Omega_{2}}|1-rz|^{-2\lambda-2}(1-|z|^{2})^{\alpha+4}dxdy$
$\leq$ $C_{2}\|(1-rz)^{-\lambda-1}\|_{\alpha+4}^{2},$
and
$\iint_{\Delta-\Omega_{1}-\Omega_{2}}\phi_{r}(z)dxdy$ $\leq$ $C_{3}\frac{\alpha+5}{\pi}\iint_{\Delta-\Omega_{1}-\Omega_{2}}(1-|z|^{2})^{\alpha+4}dxdy$
$\leq$ $C_{3}\|1\|_{\alpha+4}^{2}=C_{3}.$
By (3), (3), (3), since $1<(\alpha+4)+2$, $2\lambda+2<(\alpha+3)+2<(\alpha+4)+2$, using again the third case of Lemma 2.3, we conclude that for some constant $M>0$ it holds that
$\iint_{\Delta}\phi_{r}(z)dxdy=\Big(\iint_{\Omega_{1}}+\iint_{\Omega_{2}}+\iint_{\Delta-\Omega_{1}-\Omega_{2}}\Big)\phi_{r}(z)dxdy\leq M,$
for all $r\in(0,1)$. The claim is proved. ∎
We continue with the proof of Proposition 3.7. We only prove the case when $R$ belongs to the class $\mathcal{Q}_{III}$. As will be seen below, all other cases are covered by this proof. For a complex number $z$, we next will use $\arg z$ to denote a argument of $z$ with $\arg z\in(-\pi,\pi]$.
When $R$ belongs to the class $\mathcal{Q}_{III}$, let $\alpha+2<2\lambda<\alpha+3$. From Lemma 2.3, we get that
(3.18) $\mathbb{M}_{R}^{2}(\alpha)$ $\geq$ $\sup\limits_{r\in(0,1)}\|S(R)(re^{i\arg z_{1}}z)\|_{{\bf M}_{\alpha,\alpha+4}}^{2}$
$\geq$ $\sup\limits_{r\in(0,1)}\frac{\|S(R)(re^{i\arg z_{1}}z)(1-rz)^{-\lambda}\|_{\alpha+4}^{2}}{\|(1-rz)^{-\lambda}\|_{\alpha}^{2}}$
We establish an estimate for $\|S(R)(re^{i\arg z_{1}}z)(1-rz)^{-\lambda}\|_{\alpha+4}^{2}$.
We write ${\bf\Pi}(z):=(1-rz)^{-\lambda}$ and obtain that
$S(R)(re^{i\arg z_{1}}z){\bf\Pi}(z)=-\frac{3}{2}\frac{z_{1}^{-2}{\bf\Pi}(z)}{(1-rz)^{2}}-\frac{{z_{1}}^{-1}D_{1}(z){\bf\Pi}(z)}{1-rz}$
$-\frac{3}{2}\sum_{r=2}^{s}\frac{z_{r}^{-2}{\bf\Pi}(z)}{(1-re^{i(\arg z_{1}-\arg z_{r})}z)^{2}}+\sum_{r=2}^{s}\frac{-z_{r}^{-1}D_{r}(z){\bf\Pi}(z)}{1-re^{i(\arg z_{1}-\arg z_{r})}z}$
$-\frac{3}{2}\sum_{j=1}^{l}\frac{e^{-2i\theta_{j}}{\bf\Pi}(z)}{(1-re^{i(\arg z_{1}-\theta_{j})}z)^{2}}+\sum_{j=1}^{l}\frac{-e^{-i\theta_{j}}E_{j}(z){\bf\Pi}(z)}{1-re^{i(\arg z_{1}-\theta_{j})}z}$
$-2\sum_{k=1}^{t}\frac{e^{-2i\widetilde{\theta}_{k}}{\bf\Pi}(z)}{(1-re^{i(\arg z_{1}-\widetilde{\theta}_{k})}z)^{2}}+\sum_{k=1}^{t}\frac{-e^{-i\widetilde{\theta}_{k}}F_{k}(z){\bf\Pi}(z)}{1-re^{i(\arg z_{1}-\widetilde{\theta}_{k})}z}$
$\quad+{\bf Anal}(re^{i\arg z_{1}}z){\bf\Pi}(z).$
Here, note that $\arg z_{1}-\arg z_{r},\arg z_{1}-\theta_{j},\arg z_{1}-\widetilde{\theta}_{k}$ are pairwise distinct for all $r,j,k$. We regard the right side of (3) as a sum of the $(2s+2l+2t+1)$ terms listed above. Consequently, in the integral $\|S(R)(re^{i\arg z_{1}}z)(1-rz)^{-\lambda}\|_{\alpha+4}^{2}$, we first use the following
$\Big|\sum_{m=1}^{n}a_{m}\Big|^{2}\geq|a_{1}|^{2}-2|a_{1}|\sum_{m=2}^{n}|a_{m}|,$
with respect to (3) by taking $n=2s+2l+2t+1$ and
$a_{1}=-\frac{3}{2}\frac{z_{1}^{-2}{\bf\Pi}(z)}{(1-rz)^{2}},$
and then integrate both sides of the inequality to obtain that
$\|S(R)(re^{i\arg z_{1}}z)(1-rz)^{-\lambda}\|_{\alpha+4}^{2}\geq\|-\frac{3}{2}\frac{z_{1}^{-2}{\bf\Pi}(z)}{(1-rz)^{2}}\|_{\alpha+4}^{2}-2\bf\Sigma.$
Here, $\bf\Sigma$ denotes a sum of $2s+2l+2t$ integrals, each of which is of one of the types ${\bf T}_{1},{\bf T}_{2},{\bf T}_{3},{\bf T}_{4}$ appearing in Claim 3.8. Therefore, from Claim 3.8, we see that there is a constant ${\bf M}>0$ such that
$\|S(R)(re^{i\arg z_{1}}z)(1-rz)^{-\lambda}\|_{\alpha+4}^{2}\geq\|-\frac{3}{2}\frac{z_{1}^{-2}{\bf\Pi}(z)}{(1-rz)^{2}}\|_{\alpha+4}^{2}-{\bf M},$
for all $r\in(0,1)$. It follows from (3.18) that
(3.20) $\mathbb{M}_{R}^{2}(\alpha)$ $\geq$ $\sup\limits_{r\in(0,1)}\frac{\|-\frac{3}{2}\frac{z_{1}^{-2}{\bf\Pi}(z)}{(1-rz)^{2}}\|_{\alpha+4}^{2}-{\bf M}}{\|(1-rz)^{-\lambda}\|_{\alpha}^{2}}$
$\geq$ $\sup\limits_{r\in(0,1)}\Big(\frac{9}{4}\frac{\|(1-rz)^{-2\lambda-2}\|_{\alpha+4}^{2}}{\|(1-rz)^{-\lambda}\|_{\alpha}^{2}}-\frac{{\bf M}}{\|(1-rz)^{-\lambda}\|_{\alpha}^{2}}\Big).$
On the other hand, when $\alpha+2<2\lambda<\alpha+3$, by Lemma 2.3, we have
(3.21) $\|(1-rz)^{-\lambda}\|_{\alpha}^{2}=\frac{\Gamma(\alpha+2)\Gamma(2\lambda-\alpha-2)}{[\Gamma(\lambda)]^{2}}\cdot\frac{1+{o}(1)}{(1-r^{2})^{2\lambda-\alpha-2}},\,{\textup{as}}\,\,r\rightarrow 1^{-},$
and
(3.22) $\|(1-rz)^{-\lambda-2}\|_{\alpha+4}^{2}=\frac{\Gamma(\alpha+6)\Gamma(2\lambda-\alpha-2)}{[\Gamma(\lambda+2)]^{2}}\cdot\frac{1+{o}(1)}{(1-r^{2})^{2\lambda-\alpha-2}},\,{\textup{as}}\,\,r\rightarrow 1^{-}.$
Combining (3.20), (3.21) and (3.22), let $r\to 1^{-}$, we obtain that
$\mathbb{M}_{R}^{2}(\alpha)\geq\frac{9}{4}\frac{\Gamma(\alpha+6)}{\Gamma(\alpha+2)}\frac{[\Gamma(\lambda)]^{2}}{[\Gamma(\lambda+2)]^{2}}=\frac{9}{4}\frac{(\alpha+5)(\alpha+4)(\alpha+3)(\alpha+2)}{[\lambda(\lambda+1)]^{2}}.$
Finally, let $\lambda\to(\frac{\alpha}{2}+1)^{+}$, we have
$\mathbb{M}_{R}^{2}(\alpha)\geq\frac{9}{4}\frac{(\alpha+5)(\alpha+4)(\alpha+3)(\alpha+2)}{[(\frac{\alpha}{2}+1)(\frac{\alpha}{2}+2)]^{2}}=\frac{36(\alpha+5)(\alpha+3)}{(\alpha+2)(\alpha+4)}=\mathbb{M}_{\kappa}^{2}(\alpha).$
This completes the proof of Proposition 3.7 and hence of Theorem 1.5. ∎
4. Proof of Theorem 1.8
We first prove that
(4.1) $\mathbb{M}_{g_{0}}^{2}(\alpha,\beta)\geq\frac{1}{2^{\beta-\alpha}}\frac{\Gamma(\beta+2)}{\Gamma(\alpha+2)}\Big[\frac{\Gamma(1+\alpha/2)}{\Gamma(\beta/2+1)}\Big]^{2}.$
From Lemma 2.3, for $r\in(0,1)$, $2\lambda>\alpha+2$, we know that
$\|{(1-rz^{2})^{-\lambda}}\|_{\alpha}^{2}=\frac{\Gamma(\alpha+2)\Gamma(2\lambda-\alpha-2)}{2^{\alpha+1}[\Gamma(\lambda)]^{2}}\cdot\frac{1+{o}(1)}{(1-r^{2})^{2\lambda-\alpha-2}},\,{\textup{as}}\,\,r\rightarrow 1^{-},$
and
$\|{(1-rz^{2})^{-\lambda-\frac{\beta-\alpha}{2}}}\|_{\beta}^{2}=\frac{\Gamma(\beta+2)\Gamma(2\lambda-\alpha-2)}{2^{\beta+1}[\Gamma(\lambda+\frac{\beta-\alpha}{2})]^{2}}\cdot\frac{1+{o}(1)}{(1-r^{2})^{2\lambda-\alpha-2}},\,{\textup{as}}\,\,r\rightarrow 1^{-}.$
It follows from Lemma 2.4 that
$\mathbb{M}_{g_{0}}^{2}(\alpha,\beta)$ $\geq$ $\mathbb{M}_{g_{0}(rz)}^{2}(\alpha,\beta)\geq\sup\limits_{r\in(0,1)}\frac{\|{(1-rz^{2})^{-\lambda-\frac{\beta-\alpha}{2}}}\|_{\beta}^{2}}{\|{(1-rz^{2})^{-\lambda}}\|_{\alpha}^{2}}$
$\geq$ $\frac{1}{2^{\beta-\alpha}}\frac{\Gamma(\beta+2)}{\Gamma(\alpha+2)}\frac{[\Gamma(\lambda)]^{2}}{[\Gamma(\lambda+\frac{\beta-\alpha}{2})]^{2}}.$
Then, let $\lambda\to(\frac{\alpha}{2}+1)^{+}$, we obtain (4.1).
Next, we prove
$\mathbb{M}_{g_{0}}^{2}(\alpha,\beta)\leq\frac{1}{2^{\beta-\alpha}}\frac{\Gamma(\beta+2)}{\Gamma(\alpha+2)}\Big[\frac{\Gamma(1+\alpha/2)}{\Gamma(\beta/2+1)}\Big]^{2}.$
That is to say, for any $\phi\in\mathbf{A}_{\alpha}^{2}$, we need show that
(4.2) $\|g_{0}\phi\|_{\beta}^{2}\leq\frac{1}{2^{\beta-\alpha}}\frac{\Gamma(\beta+2)}{\Gamma(\alpha+2)}\Big[\frac{\Gamma(1+\alpha/2)}{\Gamma(\beta/2+1)}\Big]^{2}\|\phi\|_{\alpha}^{2}.$
That is
$\iint_{\Delta}(1-|z|^{2})^{\beta}|1-z|^{\beta-\alpha}|\phi(z)|^{2}dxdy$
$\leq\frac{1}{2^{\beta-\alpha}}\frac{\Gamma(\beta+1)}{\Gamma(\alpha+1)}\Big[\frac{\Gamma(1+\alpha/2)}{\Gamma(\beta/2+1)}\Big]^{2}\iint_{\Delta}|\phi(z)|^{2}(1-|z|^{2})^{\alpha}dxdy.$
We will now establish an equivalent formulation of (4.2). By using the Cayley transform
$z=z(w)=\frac{w-i}{w+i},\,\,z\in\Delta,\,\,w=u+iv\in\mathbb{C}_{+},$
which maps conformally $\mathbb{C}_{+}$ onto $\Delta$, we get that
$1-z=\frac{2i}{w+i},\,\,1-|z|^{2}=\frac{4v}{|w+i|^{2}},\,\,dxdy=\frac{4}{|w+i|^{4}}dudv.$
We set
$h(w):=\frac{\phi(z(w))}{(w+i)^{\alpha+2}}.$
A direct computation shows that the left side of (4) equals
$4^{\alpha+1}\iint_{\mathbb{C}_{+}}\frac{|h(w)|^{2}}{|w|^{\beta-\alpha}}v^{\beta}dudv,$
and the integral in the right side of (4) equals
$4^{\alpha+1}\iint_{\mathbb{C}_{+}}|h(w)|^{2}v^{\alpha}dudv.$
It follows that (4.2) is equivalent to
(4.3) $\iint_{\mathbb{C}_{+}}\frac{|h(w)|^{2}}{|w|^{\beta-\alpha}}v^{\beta}dudv\leq\frac{1}{2^{\beta-\alpha}}\frac{\Gamma(\beta+1)}{\Gamma(\alpha+1)}\Big[\frac{\Gamma(1+\alpha/2)}{\Gamma(\beta/2+1)}\Big]^{2}\iint_{\mathbb{C}_{+}}|h(w)|^{2}v^{\alpha}dudv.$
From Lemma 2.1, we know that $h\in\mathbf{A}_{\alpha}^{2}(\mathbb{C}_{+})$ so that there is a $\widetilde{h}$ that belongs to $L_{\alpha}^{2}(\mathbb{R}_{+})$ such that
$h(w)=\int_{0}^{\infty}\tilde{h}(t)e^{iwt}\,dt,\quad w\in\mathbb{C}_{+},$
and
$\|h\|_{\alpha}^{2}=\frac{\Gamma(\alpha+1)}{2^{\alpha}}\int_{0}^{\infty}|\tilde{h}(t)|^{2}\frac{dt}{t^{\alpha+1}}.$
Note that for any $v>0$, we have
$\Big(\int_{0}^{\infty}|\widetilde{h}(t)|e^{-vt}dt\Big)^{2}$ $\leq$ $\Big(\int_{0}^{\infty}|\widetilde{h}(t)|^{2}\frac{dt}{t^{\alpha+1}}\Big)\Big(\int_{0}^{\infty}e^{-2vt}t^{\alpha+1}dt\Big)$
$=$ $\frac{\Gamma(\alpha+2)}{(2v)^{\alpha+2}}\|\widetilde{h}\|_{\alpha}^{2}<\infty.$
Hence, we may define the Laplace transform of $\widetilde{h}$ by
$\mathfrak{L}(\widetilde{h})(s):=\int_{0}^{\infty}\widetilde{h}(t)e^{-st}dt,\,\,\Re(s)>0.$
Consequently,
$h(w)=\mathfrak{L}(\widetilde{h})(-iw).$
Now set $\gamma=(\beta-\alpha)/2$ and define
$H(t):=\frac{1}{\Gamma(\gamma)}\int_{0}^{t}(t-x)^{\gamma-1}\widetilde{h}(x)\,dx,\,\,t>0,$
which is a convolution and can be written as
$H(t)=\frac{1}{\Gamma(\gamma)}(k_{\gamma}*\widetilde{h})(t),\qquad k_{\gamma}(t):={t}^{\gamma-1},t>0.$
It is well known that for $\Re(s)>0$,
$\mathfrak{L}\{k_{\gamma}\}(s)=\Gamma(\gamma)\,s^{-\gamma}.$
Here, $s^{-\gamma}=e^{-\gamma\ln s}$ and $\ln s$ is a branch of the logarithmic function with $\ln 1=0$. It follows from the convolution theorem of Laplace transform that
$\mathfrak{L}(H)(s)=\mathfrak{L}(k_{\gamma})(s)\cdot\mathfrak{L}(\widetilde{h})(s)=s^{-\gamma}\,\mathfrak{L}(\widetilde{h})(s),\quad\Re(s)>0.$
Then, setting s=iws=-iw for $w\in\mathbb{C}_{+}$, since $\Re(s)=\Im(w)>0$, we can define
${\bf{H}}(w):=\int_{0}^{\infty}H(t)e^{iwt}\,dt.$
and obtain that ${\bf{H}}(w)=\mathfrak{L}(H)(-iw)$. Consequently, we have
${\bf H}(w)=(-iw)^{-\gamma}\,\mathfrak{L}(\widetilde{h})(-iw)=(-iw)^{-\gamma}h(w),\,\,w\in\mathbb{C}_{+}.$
It follows that
$|{\bf H}(w)|=\frac{|h(w)|}{|w|^{\frac{\beta-\alpha}{2}}},\,\,w\in\mathbb{C}_{+},$
and $\bf H$ belongs to ${\bf A}_{\beta}^{2}$. Hence we see from Lemma 2.1 that
$\|{\bf H}\|_{\beta}^{2}=\frac{\Gamma(\beta+1)}{2^{\beta}}\int_{0}^{\infty}|H(t)|^{2}\frac{dt}{t^{\beta+1}}.$
Therefore, we see that the inequality (4.3) is equivalent to the following
$\frac{\Gamma(\beta+1)}{2^{\beta}}\int_{0}^{\infty}|H(t)|^{2}\frac{dt}{t^{\beta+1}}$
$\leq\frac{1}{2^{\beta-\alpha}}\frac{\Gamma(\beta+1)}{\Gamma(\alpha+1)}\Big[\frac{\Gamma(1+\alpha/2)}{\Gamma(\beta/2+1)}\Big]^{2}\frac{\Gamma(\alpha+1)}{2^{\alpha}}\int_{0}^{\infty}|\tilde{h}(t)|^{2}\frac{dt}{t^{\alpha+1}}.$
Furthermore, (4.2) is equivalent to
(4.4) $\int_{0}^{\infty}\Big|\frac{1}{\Gamma((\beta-\alpha)/2)}\int_{0}^{t}(t-x)^{\frac{\beta-\alpha}{2}-1}\widetilde{h}(x)dx\Big|^{2}\frac{dt}{t^{\beta+1}}$
$\leq\Big[\frac{\Gamma(1+\alpha/2)}{\Gamma(\beta/2+1)}\Big]^{2}\int_{0}^{\infty}|\tilde{h}(t)|^{2}\frac{dt}{t^{\alpha+1}}.$
Finally, we recall the weighted Hardy inequality in Lemma 2.8. Then, the desired inequality (4.4) now follows by applying (2.19) with $r=\frac{\beta-\alpha}{2}$, $a=-\alpha-1$, and $p=2$. This proves (4.2) is true.
Remark 4.1.
We end the paper with the following
Question 4.2.
Under the assumptions and notation of Theorem 1.8, is it true that
$\mathbb{M}(\alpha,\beta)=\mathbb{M}_{g_{0}}(\alpha,\beta),$
for some $\alpha,\beta$?